### v1

@@ Abstract
Segmentation is an important analysis task for biomedical images, enabling the study of individual organelles, cells or organs. Deep learning has massively improved segmentation methods, but challenges remain in generalization to new conditions, requiring costly data annotation. Vision foundation models, such as Segment Anything Model (SAM), address this issue through broad segmentation capabilities. However, these models still require finetuning on annotated data, although with less annotations, to achieve optimal results for new conditions. As a downside, they require more computational resources. This makes parameter-efficient finetuning (PEFT) relevant for their application. We contribute the first comprehensive study of PEFT for SAM applied to biomedical segmentation by evaluating 9 PEFT methods on diverse datasets. We also provide an implementation of QLoRA for vision transformers and a new approach for resource-efficient finetuning of SAM.
@@ 1 Introduction
Segmentation is a fundamental analysis task for biomedical images. It enables the study of individual objects, such as cells and organelles in microscopy, or organs and lesions in medical imaging. Deep learning has massively advanced the field. However, the large diversity of modalities and tasks so far required different methods for specific applications, such as CellPose [Pachitariu and Stringer(2022)] and Stardist [Schmidt et al.(2018)Schmidt, Weigert, Broaddus, and Myers] for cell and nucleus segmentation, or nnU-Net [Isensee et al.(2020)Isensee, Jaeger, Kohl, Petersen, and Maier-Hein] and TotalSegmentator [Wasserthal et al.(2023)Wasserthal, Breit, Meyer, Pradella, Hinck, Sauter, Heye, Boll, Cyriac, Yang, et al.] for segmentation in radiology. Adapting any of these models to a new modality or a new segmentation task requires further data annotation for training due to limited generalisation. Annotations can only be provided by experts and is time-consuming, making adaptation costly and preventing the wider use of automatic segmentation.
@@
Vision foundation models for image segmentation, e.g. Segment Anything Model (SAM) [Kirillov et al.(2023)Kirillov, Mintun, Ravi, Mao, Rolland, Gustafson, Xiao, Whitehead, Berg, Lo, Dollár, and Girshick] or SEEM [Zou et al.(2023)Zou, Yang, Zhang, Li, Li, Wang, Wang, Gao, and Lee], promise a more unified solution. These models have been trained on large annotated datasets and address both interactive and automatic segmentation. They were recently adapted to biomedical images, resulting in foundation models for microscopy [Archit et al.(2023)Archit, Nair, Khalid, Hilt, Rajashekar, Freitag, Gupta, Dengel, Ahmed, and Pape, Israel et al.(2024)Israel, Marks, Dilip, Li, Yu, Laubscher, Li, Schwartz, Pradhan, Ates, et al.] and medical imaging [Ma et al.(2024)Ma, He, Li, Han, You, and Wang, Zhao et al.(2024)Zhao, Gu, Yang, Usuyama, Lee, Kiblawi, Naumann, Gao, Crabtree, Abel, et al., Archit et al.(2025)Archit, Freckmann, and Pape]. These models are applicable to many tasks in their respective domain without adaptation, but specific finetuning can further improve them [Archit et al.(2023)Archit, Nair, Khalid, Hilt, Rajashekar, Freitag, Gupta, Dengel, Ahmed, and Pape]. Notably, they need fewer annotated data compared to other models, requiring as few as a single labeled image [Zhou et al.(2024)Zhou, Du, and Xu].
@@
Most vision foundation models use a vision transformer (ViT) [Dosovitskiy et al.(2021)Dosovitskiy, Beyer, Kolesnikov, Weissenborn, Zhai, Unterthiner, Dehghani, Minderer, Heigold, Gelly, Uszkoreit, and Houlsby] as encoder. Hence, they typically have more parameters than previous architectures, making training more resource demanding and requiring a high-end GPU, even for small training sets. To enable efficient adaptation, parameter-efficient finetuning (PEFT) has emerged, for example through low rank adaptation of attention layers (LoRA) [Hu et al.(2021)Hu, Shen, Wallis, Allen-Zhu, Li, Wang, Wang, and Chen]. Instead of updating all model parameters, these methods either update only a small subset of parameters, or they introduce a few new parameters that are updated, while freezing the rest of the model. While PEFT has been extensively studied for large language models [Pu et al.(2023)Pu, Jain, Yin, and Kaplan, Xu et al.(2023)Xu, Xie, Qin, Tao, and Wang, Balne et al.(2024)Balne, Bhaduri, Roy, Jain, and Chadha], its application in computer vision, particularly for segmentation tasks, is less explored. Several approaches that adapt foundation models to biomedical segmentation use PEFT, e.g. [Archit et al.(2023)Archit, Nair, Khalid, Hilt, Rajashekar, Freitag, Gupta, Dengel, Ahmed, and Pape, Zhou et al.(2024)Zhou, Du, and Xu] for microscopy and [Gu et al.(2024)Gu, Dong, Yang, and Mazurowski, Zhang and Liu(2023), Wei et al.(2024)Wei, Cao, Jin, Lu, Wang, and Zhang] for medical imaging. However, they primarily use LoRA without investigating hyperparameters or alternatives. Prior work has studied PEFT for classification in natural images [Xin et al.(2024)Xin, Luo, Zhou, Du, Liu, Fan, Li, and Du] as well as classification and text-to-image generation in medical images [Dutt et al.(2024)Dutt, Ericsson, Sanchez, Tsaftaris, and Hospedales].
@@
We address this gap by studying PEFT for biomedical segmentation; for SAM and its variants $\mu$SAM [Archit et al.(2023)Archit, Nair, Khalid, Hilt, Rajashekar, Freitag, Gupta, Dengel, Ahmed, and Pape] and MedicoSAM [Archit et al.(2025)Archit, Freckmann, and Pape]. We contribute:
  • An evaluation of 9 PEFT methods on 6 microscopy and 6 medical imaging datasets.
@@
  • An implementation of quantized LoRA (QLoRA) [Dettmers et al.(2023)Dettmers, Pagnoni, Holtzman, and Zettlemoyer] for ViTs.
  • Recommendations for the use of PEFT and a workflow for efficient adaptation of SAM to new segmentation tasks.
@@
Our workflow and improvements from efficient adaptation are shown in Fig. LABEL:fig:peft_sam. We believe that our study will facilitate the use of foundation models for biomedical image segmentation. Further, it will inform future developments of PEFT for computer vision. Our code is available at https://github.com/computational-cell-analytics/peft-sam.
@@
Figure 1: Overview of PEFT-SAM. a) We study PEFT of domain-specific SAMs to enable interactive correction and efficient finetuning for improved segmentation. b) Segmentation results for three datasets from $\mu$SAM, $\mu$SAM finetuned on two images with LoRA and CellSeg1 [Zhou et al.(2024)Zhou, Du, and Xu].
@@ 2.1 Additive and Selective Finetuning
Figure 2: The original architecture of SAM, comprising image encoder, mask decoder, and prompt encoder. The image encoder is fine-tuned through additive PEFT, which introduces a few additional parameters, or selective PEFT, which updates only selected parameters. We study 9 methods; freezing the image encoder not shown.
We distinguish two types of methods: additive and selective PEFT. Selective methods update a subset of the model parameters, while freezing the rest. A simple example is freezing parts of the model, for example SAM’s image encoder, mask decoder, or prompt encoder. Freezing the image encoder, which contains most parameters, yields the greatest efficiency gain. [Archit et al.(2023)Archit, Nair, Khalid, Hilt, Rajashekar, Freitag, Gupta, Dengel, Ahmed, and Pape] have shown that this approach does not have a large negative impact on segmentation quality after finetuning. We refer to this approach as Freeze Encoder. The other selective PEFT methods we study are LayerNorm Tuning (LN Tune) [Basu et al.(2023)Basu, Massiceti, Hu, and Feizi], Bias Tuning (Bias Tune) [Cai et al.(2021)Cai, Gan, Zhu, and Han], and Attention Tuning (Attn Tune). [Touvron et al.(2022)Touvron, Cord, El-Nouby, Verbeek, and Jégou], which update only the parameters of the normalization layers, the biases, or the attention weights, respectively.
@@
Fig. LABEL:fig:architecture shows an overview of these 9 PEFT methods. We apply them only to the image encoder; prompt encoder and mask decoder are always updated. We also compare these methods with normal parameter updates (Full FT). We use SAM [Kirillov et al.(2023)Kirillov, Mintun, Ravi, Mao, Rolland, Gustafson, Xiao, Whitehead, Berg, Lo, Dollár, and Girshick] and two domain-specific models, $\mu$SAM [Archit et al.(2023)Archit, Nair, Khalid, Hilt, Rajashekar, Freitag, Gupta, Dengel, Ahmed, and Pape] for microscopy, and MedicoSAM [Archit et al.(2025)Archit, Freckmann, and Pape] for medical imaging, to initialize weights. The latter two models have finetuned SAM on a large annotated dataset from the respective domain.
@@ 2.2 Interactive Segmentation
SAM is trained on a large dataset with annotations using an objective that simulates interactive segmentation. In each iteration, this objective first samples prompts from an annotated mask, predicts the object mask, and then iteratively corrects the prediction with additional prompts sampled from the annotation. Predictions and annotations are compared with a loss function for each iteration, and the average loss is used to update parameters. We use the implementation of this procedure from [Archit et al.(2023)Archit, Nair, Khalid, Hilt, Rajashekar, Freitag, Gupta, Dengel, Ahmed, and Pape].
To evaluate interactive segmentation, we automatically derive prompts from annotated masks to segment the object and then iteratively correct the segmentation, similar to the training objective. We perform 7 correction iterations. We compute the segmentation metric (App. A) between annotations and predictions for the initial prompt and corrections. We report the results for an initial point prompt, an initial box prompt, and the last correction iteration, when starting from a point ($I_{P}$), and when starting from a box ($I_{B}$).
@@ 2.3 Automatic Instance Segmentation
During inference, the segmentation is computed by deriving seeds from the distance predictions, a mask from foreground predictions, and applying a seeded watershed inside this mask. For evaluation, the instance segmentation is compared with annotations using the mean segmentation accuracy, see App. A.
@@ 2.4 Data
For medical imaging, we use AMD-SD [Hu et al.(2024)Hu, Gao, Gao, Luo, Yang, Xiong, Chen, Lin, Xia, Yin, Deng, Ma, and Li] with 3049 optical coherence tomograms and annotations for lesions. We use JSRT [Shiraishi et al.(2000)Shiraishi, Katsuragawa, Ikezoe, Matsumoto, Kobayashi, Komatsu, Matsui, Fujita, Kodera, and Doi] with 247 images and lung and heart annotations. We use Mice TumSeg [Jensen et al.(2024)Jensen, Clemmensen, Hansen, van Krimpen Mortensen, Christensen, Kjaer, and Ripa] with 452 micro-CT volumes and tumor annotations. We use Papila [Kovalyk et al.(2022)Kovalyk, Morales-Sánchez, Verdú-Monedero, Sellés-Navarro, Palazón-Cabanes, and Sancho-Gómez] with 488 fundus images and optic cup annotations. We use MOTUM [Gong et al.(2024)Gong, Xu, Peng, Cheng, Niu, Wiestler, Hong, and Li] with 64 brain MRI volumes and tumor annotations. We use PSFHS [Chen et al.(2024b)Chen, Bai, Ou, Lu, and Wang] with 1358 ultrasound images and fetal head and pubic sympysis annotations. For the 3D datasets (MOTUM and Mice TumSeg), we perform 2D segmentation (see above).
@@ 3.1 Comparison of PEFT methods
We evaluate the PEFT methods (Sec. 2.1), on 6 microscopy datasets (Sec. 2.4) for interactive and automatic segmentation. We use SAM and $\mu$SAM (either LM or EM model) as base models. Models are trained with a batch size of 2 and 25 objects per image, using a A100 GPU with 80GB of VRAM. We use separate train / test splits for all datasets.
@@
fig:microscopy_result Refer to caption
Figure 3: PEFT results for automatic and interactive microscopy segmentation. Methods are ordered by parameter count, SAM and $\mu$SAM are used as base models. Circles show the best three results per dataset and task. See Fig. 6 for examples.
The results are shown in Fig. 3. Finetuning improves over the base models, depending on the task and the base model. Iterative segmentation ($I_{B}$, $I_{P}$) shows the smallest improvements, other tasks improve more. The effectiveness of PEFT methods differs across datasets. Full finetuning typically improves most, LoRA is the best overall PEFT method. QLoRA shows worse performance with SAM as base model, but not for $\mu$SAM. The results are summarized in Tab. 1 (SAM) and Tab. 2 ($\mu$SAM). We perform ablation studies for LoRA, Adaptformer and FacT in App. D. We also perform experiments for the 6 medical datasets. Here, we use SAM and MedicoSAM as base models and do not evaluate AIS, otherwise using the same settings. The results are shown in Fig. 4 and summarized in Tabs 3,4. We see the same trends as for microscopy.
[…]
fig:medical_result Refer to caption
Figure 4: PEFT results for medical imaging. MedicoSAM is used as domain-specific base model and we do not evaluate AIS, otherwise same settings as for microscopy.
We also evaluate the computational efficiency of PEFT methods. Tabs. 5, 6, 7 report the parameter count, memory efficiency, and training times for ViT-B, ViT-L and ViT-H. Fig. 8 shows the training times for microscopy. Overall, PEFT methods only yield marginal efficiency gains for ViT-B and ViT-L, with the exception of Freeze Encoder, and small gains for ViT-H. As a preliminary conclusion, PEFT methods generally don’t lead to better segmentation quality than full finetuning. Their efficiency gains are minor, except for Freeze Encoder. LoRA and, for small domain gaps, QLoRA provide relatively good quality-efficiency trade-offs.

### v2

@@ Abstract
Segmentation is an important analysis task for biomedical images, enabling the study of individual organelles, cells or organs. Deep learning has massively improved segmentation methods, but challenges remain in generalization to new conditions, requiring costly data annotation. Vision foundation models, such as Segment Anything Model (SAM), address this issue through improved generalization. However, these models still require finetuning on annotated data, although with less annotations, to achieve optimal results for new conditions. As a downside, they require more computational resources. This makes parameter-efficient finetuning (PEFT) relevant. We contribute the first comprehensive study of PEFT for SAM applied to biomedical images. We find that the placement of PEFT layers is more important for efficiency than the type of layer for vision transformers and we provide a recipe for resource-efficient finetuning.
@@ 1 Introduction
Segmentation is a fundamental analysis task for biomedical images. It enables the study of individual objects, such as cells and organelles in microscopy, or organs and lesions in medical imaging. Deep learning has massively advanced the field. However, the large diversity of modalities and tasks so far required different methods for specific applications, such as CellPose [Pachitariu and Stringer(2022)] and Stardist [Schmidt et al.(2018)Schmidt, Weigert, Broaddus, and Myers] for cell and nucleus segmentation, or nnU-Net [Isensee et al.(2020)Isensee, Jaeger, Kohl, Petersen, and Maier-Hein] and TotalSegmentator [Wasserthal et al.(2023)Wasserthal, Breit, Meyer, Pradella, Hinck, Sauter, Heye, Boll, Cyriac, Yang, et al.] for segmentation in radiology. Adapting any of these models to a new modality or a new task requires further data annotation for training due to limited generalization. Annotations can only be provided by experts and are time-consuming, making adaptation costly and preventing the wider use of automatic segmentation.
@@
Vision foundation models for image segmentation, e.g. Segment Anything Model (SAM) [Kirillov et al.(2023)Kirillov, Mintun, Ravi, Mao, Rolland, Gustafson, Xiao, Whitehead, Berg, Lo, Dollar, and Girshick] or SEEM [Zou et al.(2023)Zou, Yang, Zhang, Li, Li, Wang, Wang, Gao, and Lee], promise a more unified solution. These models have been trained on large annotated datasets and address both interactive and automatic segmentation. They were recently adapted to biomedical images, resulting in foundation models for microscopy [Archit et al.(2025a)Archit, Freckmann, Nair, Khalid, Hilt, Rajashekar, Freitag, Teuber, Buckley, von Haaren, Gupta, Dengel, Ahmed, and Pape, Israel et al.(2024)Israel, Marks, Dilip, Li, Yu, Laubscher, Li, Schwartz, Pradhan, Ates, et al.] and medical imaging [Ma et al.(2024)Ma, He, Li, Han, You, and Wang, Zhao et al.(2024)Zhao, Gu, Yang, Usuyama, Lee, Kiblawi, Naumann, Gao, Crabtree, Abel, et al., Archit et al.(2025b)Archit, Freckmann, and Pape]. These models are applicable to many tasks in their respective domain without adaptation, but specific finetuning can further improve them [Archit et al.(2025a)Archit, Freckmann, Nair, Khalid, Hilt, Rajashekar, Freitag, Teuber, Buckley, von Haaren, Gupta, Dengel, Ahmed, and Pape]. Notably, they need fewer annotated data compared to other models, requiring as few as a single labeled image [Zhou et al.(2024)Zhou, Du, and Xu].
@@
Most vision foundation models use a vision transformer (ViT) [Dosovitskiy et al.(2021)Dosovitskiy, Beyer, Kolesnikov, Weissenborn, Zhai, Unterthiner, Dehghani, Minderer, Heigold, Gelly, Uszkoreit, and Houlsby] as encoder. Hence, they typically have more parameters than previous architectures, making training more resource demanding and requiring a high-end GPU, even for small training sets. To enable efficient adaptation, parameter-efficient finetuning (PEFT) has emerged, for example through low rank adaptation of attention layers (LoRA) [Hu et al.(2022)Hu, Shen, Wallis, Allen-Zhu, Li, Wang, Wang, and Chen]. Instead of updating all model parameters, these methods either update only a small subset of parameters, or they introduce a few new parameters that are updated, while freezing the rest of the model. While PEFT has been extensively studied for large language models [Pu et al.(2023)Pu, Jain, Yin, and Kaplan, Xu et al.(2023)Xu, Xie, Qin, Tao, and Wang, Balne et al.(2024)Balne, Bhaduri, Roy, Jain, and Chadha], its application in computer vision, particularly for segmentation tasks, is less explored. Several approaches that adapt SAM to biomedical segmentation use PEFT, e.g. [Archit et al.(2025a)Archit, Freckmann, Nair, Khalid, Hilt, Rajashekar, Freitag, Teuber, Buckley, von Haaren, Gupta, Dengel, Ahmed, and Pape, Zhou et al.(2024)Zhou, Du, and Xu] for microscopy and [Gu et al.(2025)Gu, Dong, Yang, and Mazurowski, Zhang and Liu(2023), Wei et al.(2024)Wei, Cao, Jin, Lu, Wang, and Zhang, Wu et al.(2025)Wu, Wang, Yang, Kang, He, and Li] for medical imaging. However, they primarily use LoRA without investigating its hyperparameters or alternatives. To our knowledge, none of this prior work studied the practical efficiency gains through PEFT, relying on the trainable parameter count as a measure for efficiency. Other prior work has studied PEFT for classification in natural images [Xin et al.(2024)Xin, Luo, Zhou, Du, Liu, Fan, Li, and Du] as well as classification and text-to-image generation in medical images [Dutt et al.(2024)Dutt, Ericsson, Sanchez, Tsaftaris, and Hospedales].
@@
We address this gap by studying PEFT for biomedical segmentation; for SAM and its variants $\mu$SAM [Archit et al.(2025a)Archit, Freckmann, Nair, Khalid, Hilt, Rajashekar, Freitag, Teuber, Buckley, von Haaren, Gupta, Dengel, Ahmed, and Pape] and MedicoSAM [Archit et al.(2025b)Archit, Freckmann, and Pape]. We contribute:
  • An evaluation of several PEFT methods on microscopy and medical imaging.
  • A thorough evaluation of efficiency gains due to PEFT.
@@
  • Implementation of more efficient “late PEFT” approaches based on these finding and an implementation of quantized LoRA (QLoRA) [Dettmers et al.(2023)Dettmers, Pagnoni, Holtzman, and Zettlemoyer] for ViTs.
  • Recommendations for the use of PEFT and a recipe for efficient adaptation of SAM.
@@
Our workflow and improvements from efficient adaptation are shown in Fig. 1. We believe that our study will facilitate the use of foundation models for biomedical image segmentation. Further, it will inform future developments of PEFT for computer vision. Our code is available at https://github.com/computational-cell-analytics/peft-sam.
@@
Figure 1: Overview of PEFT-SAM. a) We study PEFT of domain-specific SAMs to enable interactive correction and efficient finetuning for improved segmentation. b) Segmentation results for three datasets from $\mu$SAM, $\mu$SAM finetuned with LoRA and CellSeg1, all trained on two images [Zhou et al.(2024)Zhou, Du, and Xu].
@@ 2.1 Additive and Selective Finetuning
Figure 2: The original architecture of SAM, comprising image encoder, mask decoder and prompt encoder, with an additional segmentation decoder, proposed by $\mu$SAM. The image encoder is fine-tuned through additive PEFT, which introduces a few additional parameters, or selective PEFT, which updates only selected parameters. We also introduce late PEFT, which adapts only layers at the end of ViT.
We distinguish two types of methods: additive and selective PEFT. Selective methods update a subset of the model parameters, while freezing the rest. A simple example is freezing parts of the model, for example SAM’s image encoder, mask decoder, or prompt encoder. Freezing the image encoder, which contains most parameters, yields the greatest efficiency gain. [Archit et al.(2025a)Archit, Freckmann, Nair, Khalid, Hilt, Rajashekar, Freitag, Teuber, Buckley, von Haaren, Gupta, Dengel, Ahmed, and Pape] have shown that this approach does not have a large negative impact on segmentation quality after finetuning. We refer to this approach as Freeze Encoder. The other selective PEFT methods we study are LayerNorm Tuning (LN Tune) [Basu et al.(2024)Basu, Hu, Massiceti, and Feizi], Bias Tuning (Bias Tune) [Cai et al.(2020)Cai, Gan, Massachusetts Institute of Technology, and Massachusetts Institute of Technology], and Attention Tuning (Attn Tune). [Touvron et al.(2022)Touvron, Cord, El-Nouby, Verbeek, and Jégou], which update only the parameters of the normalization layers, the biases, or the attention weights, respectively.
@@
Fig. 2 shows an overview of these 9 PEFT methods. We apply them only to the image encoder; prompt encoder and mask decoder are always updated. Further, we study an approach we call “late PEFT”, where we insert trainable parameters / unfreeze weights only in the late layers of the image encoder. This is motivated by our findings on computational efficiency, see App. A. Specifically, we study “late LoRA”, “late QLoRA“ and “late finetuning”. We compare all PEFT methods with normal parameter updates (Full FT). We use SAM [Kirillov et al.(2023)Kirillov, Mintun, Ravi, Mao, Rolland, Gustafson, Xiao, Whitehead, Berg, Lo, Dollar, and Girshick] and two domain-specific models, $\mu$SAM [Archit et al.(2025a)Archit, Freckmann, Nair, Khalid, Hilt, Rajashekar, Freitag, Teuber, Buckley, von Haaren, Gupta, Dengel, Ahmed, and Pape] for microscopy, and MedicoSAM [Archit et al.(2025b)Archit, Freckmann, and Pape] for medical imaging, to initialize weights. The latter two models have finetuned SAM on a large annotated dataset for the respective domain.
@@ 2.2 Interactive Segmentation
SAM is trained on a large dataset with annotations using an objective that simulates interactive segmentation. In each iteration, this objective first samples prompts from an annotated mask, predicts the object mask, and then iteratively corrects the prediction with additional prompts sampled from the annotation. Predictions and annotations are compared with a loss function for each iteration, and the average loss is used to update parameters. We use the implementation of this procedure from [Archit et al.(2025a)Archit, Freckmann, Nair, Khalid, Hilt, Rajashekar, Freitag, Teuber, Buckley, von Haaren, Gupta, Dengel, Ahmed, and Pape].
To evaluate interactive segmentation, we automatically derive prompts from annotations to segment the object and then iteratively correct it, similar to the training objective. We perform 7 correction iterations. We compute the metric (App. D) between annotations and predictions for the initial prompt and corrections. We report the results for an initial point prompt, an initial box prompt, and the last correction iteration, when starting from a point ($I_{P}$, appendix only), and when starting from a box ($I_{B}$, appendix only).
@@ 2.3 Automatic Instance Segmentation
During inference, the segmentation is computed by deriving seeds from the distance predictions, a mask from foreground predictions, and applying a seeded watershed inside this mask. For evaluation, the instance segmentation is compared with annotations using the mean segmentation accuracy, see App. D.
@@ 2.4 Data
For medical imaging, we use AMD-SD [Hu et al.(2024)Hu, Gao, Gao, Luo, Yang, Xiong, Chen, Lin, Xia, Yin, Deng, Ma, and Li] with 3049 optical coherence tomograms and annotations for lesions. We use JSRT [Shiraishi et al.(2000)Shiraishi, Katsuragawa, Ikezoe, Matsumoto, Kobayashi, Komatsu, Matsui, Fujita, Kodera, and Doi] with 247 images and lung and heart annotations. We use Mice TumSeg [Jensen et al.(2024)Jensen, Clemmensen, Hansen, van Krimpen Mortensen, Christensen, Kjaer, and Ripa] with 452 micro-CT volumes and tumor annotations. We use Papila [Kovalyk et al.(2022)Kovalyk, Morales-Sánchez, Verdú-Monedero, Sellés-Navarro, Palazón-Cabanes, and Sancho-Gómez] with 488 fundus images and optic cup annotations. We use MOTUM [Gong et al.(2024)Gong, Xu, Peng, Cheng, Niu, Wiestler, Hong, and Li] with 64 brain MRI volumes and tumor annotations. We use PSFHS [Chen et al.(2024b)Chen, Bai, Ou, Lu, and Wang] with 1358 ultrasound images and fetal head and pubic symphysis annotations. For the 3D datasets (MOTUM and Mice TumSeg), we perform 2D segmentation (see above).
@@ 3.1 Comparison of PEFT methods
We evaluate the PEFT methods (Sec. 2.1) on 6 microscopy and 6 medical datasets (Sec. 2.4), using the complete datasets with a train / test split. We use SAM as base model in all cases, $\mu$SAM (either LM or EM model) for microscopy, and MedicoSAM for medical data. For microscopy we evaluate interactive and automatic segmentation via AIS (Sec. 2.3), for medical data we only evaluate interactive segmentation. Models are trained with a batch size of 2 and 25 objects per image, using a A100 GPU with 80GB of VRAM.
@@
fig:peft_seg_result Refer to caption
Figure 3: PEFT results for microscopy (left) and medical (right) segmentation. Methods are ordered by parameter count, SAM and $\mu$SAM / MedicoSAM are used as base models. Circles show the best three results per dataset / task. See also Fig. 8 and Fig.  9.
The results are shown in Fig. 3. Finetuning always improves, with the magnitude depending on the task / dataset and the base model. Full FT yields the most consistent improvement, with Attention Tuning, LoRA, Late LoRA and Late FT typically improving similarly. The improvements from other PEFT methods are not quite as consistent. Interestingly, (Late) QLoRA shows worse performance with SAM as base model, but not for $\mu$SAM / MedicoSAM. The results are summarized and discussed in more detail in App. E.1, including results for further segmentation tasks ($I_{B}$, $I_{P}$) and further medical datasets. We perform ablation studies for LoRA, Adaptformer and FacT in App. G.
We also evaluate the computational efficiency of PEFT methods. The detailed results on efficiency are given in App. E.2. Our most notable finding is that our tasks are constrained by the size of activations rather than trainable parameter count, see App. A. Consequently, most prior PEFT methods yield only a marginal efficiency gain, with the exception of freezing the encoder. We introduced ”late LoRA“, ”late QLoRA“, and ”late finetuning“ (late FT) due to this observation. They achieve meaningful efficiency gains while improving segmentation quality compared to full freezing of the encoder, see the respective results in Fig. 3. We analyze different late PEFT settings in detail in Fig. 5 and chose the best parameter settings based on these results for other experiments.